OpenStudy (anonymous):

the equation of the tangent to the curve y=xsinx at the point (pi/2, pi/2) is

OpenStudy (zzr0ck3r):

take the derivative = sin(x) + x*cos(x)

OpenStudy (zzr0ck3r):

now you have the function that gives you slope plug in pi/2 sin(pi/2) + (pi/2)(cos(pi/2) =1 y = mx+b but m = 1 y = x+b pi/2=pi/2+b b = 0 so y = x is the equation tangent to the curve at that point